OpenStudy (anonymous):

write an equation for the line that is parallel to the given line and that passes through the given point. y=1/2x-8;(2,-15)

OpenStudy (saifoo.khan):

Parallel means same slope. The slope for the line is, 1/2 Now you are given a slope and point (2,-15) use y - y1 = m(x - x1) and solve.

OpenStudy (callisto):

For more information, (x1, y1) = (2, -15) and m =1/2 and substitute them into the equation